Janet is presently one-sixth as old as her
mother. Her mother’s age, when divided
by 2, 3, 4, 6, or 8, always leaves one
remaining year. This is not so when her
mother’s present age is divided by 5, for
then there is nothing left over. How old
is Janet’s mother?
How old is Janet?

Found 3 solutions by Theo, ikleyn, greenestamps:
Answer by Theo(13342)

m = 25
25 = 2 * 12 + 1
25 = 3 * 8 + 1
25 = 4 * 6 + 1
25 = 6 * 4 + 1
25 = 8 * 3 + 1
25 = 5 * 5 + 0
her present age is 25.


Answer by ikleyn(52776)

.
Janet is presently one-sixth as old as her mother.
Her mother’s age, when divided by 2, 3, 4, 6, or 8,
always leaves one remaining year.
This is not so when her mother’s present age is divided by 5,
for then there is nothing left over. How old is Janet’s mother?
How old is Janet?
~~~~~~~~~~~~~~~~~~~~~~~~


As it is worded, printed and presented in this post, this problem
is posed incorrectly, is self-contradictory and, speaking mathematically,
is DEFECTIVE.


I will not explain WHY, since to explain it to a mathematically illiterate
person is counterproductive.

Mathematically illiterate person should not create/compose "math problems"
and distribute them in the Internet.


This my note is addressed to other visitors, who probably will read this post.
Simply ignore it and do not consider this "problem" seriously and do not spend your time for nothing.


The fact that @Theo presented some solution, means NOTHING.

In order for the problem be correctly posed, its condition should be changed/re-written.



Answer by greenestamps(13198)



We can guess what the proper presentation of the problem is supposed to be.... But as presented, the problem is nonsense.

Word problems involving people's ages only make sense if the ages are whole numbers of years.

In this problem, the mother's age divided by 2 leaves 1 year left over, so the mother's age in years is an odd number.

But Janet's age in years can't be one-sixth of an odd number, because one-sixth of an odd number is never a whole number.